OpenStudy (anonymous):

Which of the following best describes a way in which individuals have an impact on the level of greenhouse gases in the atmosphere? A. Using aerosols that contain chlorofluorocarbons (CFCs) decreases the amount of protective ozone (O3) in the atmosphere. B. Using appliances that absorb radiation decreases the amount of nitrous oxide (N2O) in the atmosphere. C. Burning fossil fuels decreases the amount of water vapor (H2O) in the atmosphere. D. Deforestation decreases the amount of carbon dioxide (CO2) in the atmosphere

OpenStudy (anonymous):

@Abhisar

OpenStudy (abhisar):

What do you think ?

OpenStudy (anonymous):

D

OpenStudy (abhisar):

Plants use carbon-dioxide to make glucose. So deforestation i.e cutting down of trees increases the carbon dioxide level.

OpenStudy (abhisar):

What do you think about option A ?

OpenStudy (abhisar):

????

OpenStudy (anonymous):

the gases mentioned in above question are all grennhouse gases .. O3,N2O,CO2 and water vapours too.. decrease or increase in any1 of these wil affect greaanhouse... By their percentage contribution to the greenhouse effect on Earth the four major gases are: water vapor, 36–70% carbon dioxide, 9–26% methane, 4–9% ozone, 3–7% a should C the answer if we r talkiing about impact on green house effect.. but question is asked which factor wil hav impact on level of "greenhouse gases" each factor is affecting level of greenhouse gases ... i think there is a mistake in question or we need to study more .. :) @Nicksl99 can u check ur question again plz is it ryt?

OpenStudy (anonymous):

@Abhisar hw u think its A?

OpenStudy (abhisar):

Can you explain me how burning of fossil fuels decrease the amount of water vapor @saroash ?

OpenStudy (anonymous):

can u guys help me on another question. its for english

OpenStudy (anonymous):

@Abhisar i used the fact that burning furnace or heater in a room decreaes humitdity in room... but i could b wrong because it happens in a closed room

OpenStudy (anonymous):

in room relative humidity is decareaed not absolute .. :/ just read now sory it may r may not b A

OpenStudy (abhisar):

You are talking at very small scale. Here it is being talked about Atmosphere. humidity is due to the presence of tiny water droplets in the air. WHEN WE BURN FOSSIL FUELS heat is produced and these droplets get converted into gaseous form (there is a lill difference between vapor and gas).

OpenStudy (abhisar):

Also burning of fossil fuels produce large amount of carbon dioxide which is a much more potent green house gas.

OpenStudy (anonymous):

oh... yup u r ryt ... C is eliminated u gave reason .... D is wrong too b/c if deforestation wil occur less CO2 wil b consumed ,here wrong statement is written... A is right ... as O3 is is greenhouse gas, aerosal does effect it ... hmm... now i am getting it :) @Abhisar i have no idea about statement B ..can u please explain that too wil b thankful :)

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @saroash @Abhisar i used the fact that burning furnace or heater in a room decreaes humitdity in room... but i could b wrong because it happens in a closed room \(\color{blue}{\text{End of Quote}}\) This reason has nothing to do specifically with fossil fuels. How would you explain it in desert areas where humidity is already less. ?

OpenStudy (anonymous):

i am sorry for this .... i got it now sir!

OpenStudy (abhisar):

It's ok ! You don't need to be sorry :) \(\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \Large\cdot \quad \cdot\\ \cdot\\ \huge \smile \end{matrix} \right)}}\normalsize \\ \;/\quad \;\;\quad \backslash\)

OpenStudy (abhisar):

@saroash Nitrous oxide is also a green house gas. the option B doesn't makes any sense to me. The statement "Using appliances that absorb radiation decreases the amount of nitrous oxide (N2O) in the atmosphere." is false.

OpenStudy (anonymous):

hmmm .... i got it too... wel i m learner sir .. and leearners do mak mistakes to learn more..i asked u to giv reason for A .. i was trying to say u r wrong ... i was trying to see hw am i wrong where i m making mistake to understand.... u are a senior i wanted guidance nothing else sory again ... well thanks for guiding :)